OpenStudy (anonymous):

Find the last two didgits of 6 raised to the power of 13 A) 96 B) 36 C) 76 D)16

OpenStudy (anonymous):

6^13 = 13060694016 now guess your answer

OpenStudy (anonymous):

? ?

OpenStudy (zzr0ck3r):

is this number theory?

OpenStudy (zzr0ck3r):

In number theory they ask this question but we cant use technology.

OpenStudy (anonymous):

13060694016 which are the last two digits of the above number

OpenStudy (zzr0ck3r):

yes, but you used a calculator...

OpenStudy (zzr0ck3r):

there is a way to do it without.

OpenStudy (anonymous):

Sorry , I get it now :)

OpenStudy (anonymous):

Thanks guys :3

OpenStudy (anonymous):

welcome

OpenStudy (zzr0ck3r):

ok looks like this is fine:) good because I forgot how to do it with mod10

OpenStudy (anonymous):

There is a pattern to observe:$$6^1=\color{red}6\\6^2=\color{blue}3\color{red}6\\6^3=2\color{blue}1\color{red}6\\6^4=12\color{blue}9\color{red}6\\6^5=67\color{blue}7\color{red}6\\6^6=406\color{blue}5\color{red}6\\6^7=2439\color{blue}3\color{red}6\\6^8=15036\color{blue}1\color{red}6$$ The last digit is *always* 6 and the 10's digit follows a regular pattern: \(3\to1\to9\to7\to5\to3\to1\to9\to7\to5\to3\to1\to\dots\) Notice this pattern repeats a number every 5 digits; this means exponents that are a multiple \(5\) apart will have the same second digit. What about \(13\)? Well:$$6^{13}=6^{8+5}=6^{3+5+5}$$... so \(6^{13}\) ought to have the same last digits as \(6^8\) and \(6^3\), which we can see is \(16\).